The micrometer (1 μm) is often called the micron. (a) How many microns make up 2.6 km? (b) How many centimeters equal 2.6 μm? (c) How many microns are in 2.6 yd?
Answer #1

a) 1 um = 10^-6 m

2.6 km = 2600 m =

1 m = 10^6 um

= 2600*10^6 = 2.6*10^9 um

b) 1 cm = 10^-2 m

2.6 um = 2.6*10^-6 m

= 2.6*10^-6 *10^2 = 2.6*10^-4 cm

(C) 1 yd = 0.9144 m

2.6 yd = 2*0.9144 = 2.37744 m

= 2.37744*10^6 um